OpenStudy (anonymous):

Show the following pairs of lines intersect. Determine the coordinates of the point of intersection and the angles formed by this lines. a. L1: x = 5 + 2t, y =-3 + t and L2: x = 23 - 2s, y = 6 - s What I did so far is: L1: coordinates (5, -3) direction vector (2, 1) L2: coordinates (23 , 6) direction vector ( -2, -1) I'm thinking that, this two lines are collinear because direction vectors are scalar multiple and therefore the angle form will be 0. Am I right?

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

@thomaster

OpenStudy (psymon):

Yeah, the direction vector is a scalar multiple, so parallel at the very least.

OpenStudy (psymon):

Tagged me just as I logged on, haha

OpenStudy (anonymous):

ok, thanks again:) lol..of course, i know you can help me..

OpenStudy (psymon):

For something we're learning at the same time, maybe, lol.

OpenStudy (anonymous):

yes..I'm almost done with this chapter, i just need to finish all this questions.. my next topic will be "The intersection of a line with a Plane and the intersection of two lines"

OpenStudy (psymon):

Yeah, I need to keep up with ya. Well, assuming we didn't notice that they were scalar multiples, do you know how wed check?

OpenStudy (anonymous):

...no, i'm not sure, there's a question like that and i'm doing it now.. trying to figure it out..

OpenStudy (psymon):

Im just taking out my text, just got home, lol.

OpenStudy (anonymous):

oh you are attending school?

OpenStudy (psymon):

But I believe we'd have to solve a system of equations. So if the lines intersect, then there will be some t that will cause things to be equal. And yeah, Im a college student. School doesnt start till the 26th, though. This is just me doing my own calc 3 preview, lol.

OpenStudy (psymon):

x= 5 + 2t y = -3 + t x = 23-2s y = 6 - s

OpenStudy (psymon):

So we have to solve for s and t in this system by setting the x parametrics and the y parametrics equal: 5+2t = 23-2s -3+t = 6-s We good on how to solve systems of equations?

OpenStudy (anonymous):

like for example: \[L1: \frac{ x+3 }{ 3 }=\frac{ y+1 }{ 4 }\] \[L2:\frac{ x-6 }{3}=\frac{ y-2 }{ -2 }\] do you think I still need to change it to Cartesian equation?

OpenStudy (anonymous):

oopsss!, i did not know you post something

OpenStudy (psymon):

xDD

OpenStudy (anonymous):

Oh.. ok, I get it now

OpenStudy (anonymous):

do you still have any reminders that I need to know??:)

OpenStudy (psymon):

Not really. Im still learning it, too xD

OpenStudy (anonymous):

wait a sec. i'll try to solve another problem like this..

OpenStudy (psymon):

Alrighty. I need more examples anyway

OpenStudy (psymon):

Lot of typing there xD

OpenStudy (anonymous):

I'm confused again..:] for the example I gave above, I change it to parametric equation: \[L1: x=-3+3s, y=-1+4s\] \[L2: x=6 + 3t, y=2-2t\] then I followed what you said: x-values: \[-3+3s=6+3t\] \[3s - 3t =9\] \[s= t+3\]---> equation 1 y-values: \[-1+4s=2-2t\] \[4s+2t=3\] --->equation 2 I tried to solved it, using substitution: substitute equation 1 in equation 2 then if I will try to solve for s, it is impossible